1. ) You are given the following information for Gandolfino Pizza Co.: sales = $42,000; costs = $22,200; addition to retained earnings = $5,350; dividends paid = $1,800; interest expense = $4,600; tax rate = 35 percent. Calculate the depreciation expense. (Do not round intermediate calculations and round your final answer to nearest whole dollar amount.

B.) Red Hawk, Inc., is obligated to pay its creditors $7,100 during the year. What is the market value of the shareholders’ equity if assets have a market value of $8,750? What is the market value of the shareholders’ equity if assets equal $6,500?
Answer #1

Net income = (Sales - costs - depreciation - interest) * (1 - tax rate)

5350 + 1800 = (42000 - 22200 - depreciation - 4600) * (1 - 35%)

11000 = 15200 - depreciation

Depreciation = 4200

Market value of equity = 8750 - 7100 = 1650.

Market value of equity = 6500 - 7100 = 0

Since the value of equity cannot go below 0.